Albany Corporation purchased equipment on April 1 of Year 1 for $75,000. The asset does not have a residual value, and is estimated to be in service for 8 years. Calculate the depreciation for Year 1 using the double-declining-balance method. Round to the nearest dollar. Calculate the depreciation for Year 2 using the double-declining-balance method. Round to the nearest dollar.
Answer #1

Year

Depreciation expense using DDB

1

$   14,063

2

$   15,234

Working

Double declining Method

A

Cost

$ 75,000.00

B

Residual Value

$                 -  

C=A - B

Depreciable base

$ 75,000.00

D

Life [in years]

8

E=C/D

Annual SLM depreciation

$    9,375.00

F=E/C

SLM Rate

12.50%

G=F x 2

DDB Rate

25.00%

Year

Beginning Book Value

Depreciation rate

Depreciation expense

Ending Book Value

Accumulated Depreciation

1

$        75,000

25.00%

$   14,063

$    60,938

$ 14,063

2

$        60,938

25.00%

$   15,234

$    45,703

$ 29,297

In year 1 depreciation is for 9 months